OpenStudy (anonymous):

Integrate 1/((x-1)(x+2))

OpenStudy (amistre64):

it looks ready to turn into partial fractions

OpenStudy (amistre64):

\[\frac{1}{(x-1)(x+2)}=\frac{A}{x-1}+\frac{B}{x+2}\] when x=1; A = 1/3 when x = -2; B = -1/3 if i see it right

OpenStudy (amistre64):

\[\int\frac{1}{3(x-1)}-\frac{1}{3(x+2)}dx\]

OpenStudy (anonymous):

yes you do the "french method" and you see it right away.